THE PERMANENT ELECTRIC DIPOLE MOMENTS AND MAGNETIC HYPERFINE INTERACTON IN THE A^{2}Π STATE OF YTTRIUM MONOSULFIDE, YS

\underline{WILTON L. VIRGO} and TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

The optical Stark effect in the (0,0) A^{2}Π−X^{2}Σ^{+} band systems of a yttrium monosulfide, YS , supersonic molecular beam sample have been analyzed to produce permanent electric dipole moments, μ, for the A^{2}Π_{3/2}, and A^{2}Π_{1/2} states of 5.9(2) D, and 6.8(1) D, respectively. Fine structure splitting in the field free A^{2}Π_{3/2} − X^{2}Σ^{+} and A^{2}Π_{1/2} − X^{2}Σ^{+} spectra were analyzed to produce the magnetic hyperfine spectroscopic parameters a = -36(6) MHz, c = 111(7) MHz, and d = -107(6) MHz for the A^{2}Π state. Transition frequencies of the low rotational lines in the (0,0) A^{2}Π_{r} − X^{2}Σ^{+} band system were measured and analyzed to produce the first complete set of fine structure parameters for the A^{2}Π_{r} state.